OpenStudy (anonymous):

For what value of c is the curve y=c/(x+1) tangent to the line through the points (0,3) and (5,-2)?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

so i know y'=(x-c+1)/(x+1)^2 and i know the line of that tangent line has slope -1 and i was thinking i set them equal to each other but then i get x^2+x+c=0 the discrimnant is -b-4ac=1-4(1)c we want the discrimnant to be zero so we have one unique tangent line right? so we have 1-4c=0 this implies c=1/4

OpenStudy (anonymous):

but the back of the book has 4

OpenStudy (anonymous):

ok the slope of the line is -1, but i think \[y'=\frac{-c}{(x+1)^2}\]

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

now if you set = -1 i bet you get the right answer

OpenStudy (anonymous):

so that is the correct way i mentioned above?

OpenStudy (anonymous):

just set the discriminant =0?

OpenStudy (anonymous):

think so let me check

OpenStudy (anonymous):

hmmm x=-1+ or - sqrt{c} don't know how to get 4 from this

OpenStudy (anonymous):

oh wait. the curve has to be tangent to the line, so i must touch the line yes?

OpenStudy (anonymous):

the line must touch the curve yes!

OpenStudy (anonymous):

so we are trying to find the intersection so maybe i should find the line of the tangent line and set it equal to the curve y=c/(x+1)?

OpenStudy (anonymous):

so if my thinking is correct we have two equations, \[(x+1)^2=c\] and \[\frac{c}{x+1}=-x+3\]

OpenStudy (anonymous):

the first from the derivative being -1 and the second because for that value of c the line which is y = -x+3 must intersect the curve

OpenStudy (anonymous):

how did you get y=-x+3?

OpenStudy (anonymous):

so we geet \[(x+1)^2=(x+1)(-x+3)\]

OpenStudy (anonymous):

oh line w/ slope -1 through (0,3)

OpenStudy (anonymous):

oh duh nvm

OpenStudy (anonymous):

you've got the slope and the y intercept. i think this gives \[x+1=-x+3\] \[2x=2\] \[x=1\] but apparently that is not the answer so let me try again

OpenStudy (anonymous):

so i get x=1 and x=-1 from that if x=1 then x=(1+1)^2=2^2=4 but if x=-1 then x=(-1+1)^2=0^2=0

OpenStudy (anonymous):

those are suppose to be c i accidently wrote x again sorry

OpenStudy (anonymous):

x=1 c=(1+1)^2=4 x=-1 c=(0)^2=0

OpenStudy (anonymous):

oh right! i solved for x and forgot we were looking for c. duh

OpenStudy (anonymous):

gj thanks satellite remember that is my bike

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

amistre gave it to me

OpenStudy (anonymous):

yeah, but as you can see i stole it

OpenStudy (anonymous):

satellite do you know who i am?

OpenStudy (anonymous):

nope? myininaya?

OpenStudy (anonymous):

lol i didn't wanna seem like a retard asking it has the goddess myininaya

OpenStudy (anonymous):

as*

OpenStudy (anonymous):

ho ho ho ho ho ho or as we say here lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

glad it was me that caught this one!

OpenStudy (anonymous):

now i can retire for the night

OpenStudy (anonymous):

good night satellite thanks for showing me my error and i totally forgot to set those lines/curves equal you're the best

OpenStudy (anonymous):

no you are!

OpenStudy (anonymous):

no you!